the salary of a Man starts at Rs 800 and increase annually by Rs 20 until a maximum of Rs 1120 is reached .Find the total amount of money earned by him when he serves 15 years and 22 years.
Found 2 solutions by LinnW, richwmiller:
Answer by LinnW(1048)

for 15 years
800 +
800 + 20 +
800 + 20 + 20 +
800 + 20 + 20 + 20 +
800 + 20 + 20 + 20 + 20 +
800 + 20 + 20 + 20 + 20 + 20 +
800 + 20 + 20 + 20 + 20 + 20 + 20 +
800 + 20 + 20 + 20 + 20 + 20 + 20 + 20 +
800 + 20 + 20 + 20 + 20 + 20 + 20 + 20 + 20 +
800 + 20 + 20 + 20 + 20 + 20 + 20 + 20 + 20 + 20 +
800 + 20 + 20 + 20 + 20 + 20 + 20 + 20 + 20 + 20 + 20 +
800 + 20 + 20 + 20 + 20 + 20 + 20 + 20 + 20 + 20 + 20 + 20 +
800 + 20 + 20 + 20 + 20 + 20 + 20 + 20 + 20 + 20 + 20 + 20 + 20 +
800 + 20 + 20 + 20 + 20 + 20 + 20 + 20 + 20 + 20 + 20 + 20 + 20 + 20 +
800 + 20 + 20 + 20 + 20 + 20 + 20 + 20 + 20 + 20 + 20 + 20 + 20 + 20 + 20
= amount earned in 15 years
Notice that the column with 800's has 15 entries
The second column from the left has 14 20's
The third column has 13 20's
so the amount earned in 15 years is
15(800) +(14+13+12+11+10+9+8+7+6+5+4+3+2+1)20
= 15(800) +(105)20
= 12000 + 2100
= 14100
800+
800+20+
800+20+20+
800+20+20+20+
800+20+20+20+20+
800+20+20+20+20+20+
800+20+20+20+20+20+20+
800+20+20+20+20+20+20+20+
800+20+20+20+20+20+20+20+20+
800+20+20+20+20+20+20+20+20+20+
800+20+20+20+20+20+20+20+20+20+20+
800+20+20+20+20+20+20+20+20+20+20+20+
800+20+20+20+20+20+20+20+20+20+20+20+20+
800+20+20+20+20+20+20+20+20+20+20+20+20+20+
800+20+20+20+20+20+20+20+20+20+20+20+20+20+20+
800+20+20+20+20+20+20+20+20+20+20+20+20+20+20+20+
800+20+20+20+20+20+20+20+20+20+20+20+20+20+20+20+20+ = 1120
800+20+20+20+20+20+20+20+20+20+20+20+20+20+20+20+20+
800+20+20+20+20+20+20+20+20+20+20+20+20+20+20+20+20+
800+20+20+20+20+20+20+20+20+20+20+20+20+20+20+20+20+
800+20+20+20+20+20+20+20+20+20+20+20+20+20+20+20+20+
800+20+20+20+20+20+20+20+20+20+20+20+20+20+20+20+20
Notice that the column with 800's has 22 entries
The second column from the left has 21 20's
The third column has 20 20's
so the amount earned in 22 years is
22(800) +(21+20+19+18+17+16+15+14+13+12+11+10+9+8+7+6)20
Notice that years 18 through 22 are topped at 1120
= 22(800) +(216)20
= 17600 + 4320
= 21920 for 22 years

Answer by richwmiller(17219)